# Principal component analysis balancing prediction and approximation accuracy for spatial data

### v1

@@ 21      Calculate the $l$th PC score: $\tilde{u}_{l}\leftarrow Y^{(l)}\tilde{v}_{l}$;
Theorem 1, which is proved in Appendix A, establishes the optimality of $\tilde{U}$ and $\tilde{V}$. Numerical examples in Appendix B.1 empirically verify the optimality.
@@
For each PC (indexed by $l$), denote the singular value decomposition of $Y^{(l)}$ as $Y^{(l)}=\mathcal{S}^{(l)}{\mathcal{D}^{(l)}}{\mathcal{T}^{(l)}}^{\top}$. Take $Z,P,q$ and $\tilde{v}_{l}$ to be the specified form in Algorithm 1. Then $(\tilde{v}_{l},\tilde{\alpha}_{l},\tilde{\beta}_{l})$ is the global minimizer of (3), where $\left[\tilde{\alpha}_{l}^{\top}\quad\sqrt{\frac{\lambda_{2}}{\lambda_{1}}}\tilde{\beta}_{l}^{\top}\right]^{\top}=(\gamma Z^{\top}Z+\lambda_{1}P)^{-1}(\gamma Z^{\top}\mathcal{S}^{(l)}{\mathcal{D}^{(l)}}q)$.
The computational complexity of the procedure described above is $O\left((n+m)^{3}+n^{2}p+np^{2}\right)$, where we recall that $n$ is the sample size, $p$ is the dimensionality of $Y$, and $m$ is the number of spline basis functions included in $B$. This algorithm is computationally simple in that it obtains the optimal solution with an explicit expression, as opposed to iterative numerical optimization procedures. This algorithm depends on the specification of several quantities. For the hyper-parameters $\gamma,\lambda_{1},\lambda_{2}$, one could adopt cross-validation and choose the combination that minimizes TMSE or other metrics (e.g., MSPE/MSRE) depending on the targeted use case. Taking TMSE as an example, for the $l$th PC and for each combination of $\gamma,\lambda_{1},\lambda_{2}$, we calculate $\tilde{u}_{l},\tilde{v}_{l}$ on the training set, predict $\hat{u}_{\text{tst}}$ on the validation set, and then examine $\text{TMSE}_{l}=n^{-1}\left\lVert Y^{(l)}_{\text{tst}}-\hat{u}_{\text{tst}}\tilde{v}_{l}^{\top}\right\rVert_{2}^{2}$ to choose the optimal combination. The dimension $m$ of $B$ is not critical due to the inclusion of penalty on $\beta$ (Wood, 2017). One can start with a large $m$ (e.g., close to the sample size $n$) and the degree of freedom of $\beta$ will be controlled data-adaptively via $\lambda_{2}$.
@@ 4 Simulations
In this section, we illustrate the performance of RapPCA with simulated data, in comparison to classical and predictive PCA. In particular, we investigate three scenarios with different trade-offs between predictability and representability resulting from different data generating mechanisms.
@@
$\text{PC}^{(l)}=f_{l}(X)+\epsilon_{l},\ l=1,\ldots,6,$
@@
$Y_{n\times p}=\text{PC}_{n\times 6}M_{6\times p}+\varepsilon_{n\times p}$
where $M$ is adjusted differently in each scenario to control the contribution from different (e.g., predictable versus unpredictable) PCs, and the entries of the noise terms $\varepsilon$ are drawn from independent Normal$(0,0.1)$ distributions.
We examine the metrics of interest including TMSE, MSPE and MSRE-trn for 100 replicates of data, in terms of their overall magnitudes as well as the breakdown by each PC and/or $\gamma$ (the tuning parameter). The optimal combination of $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ is selected by minimizing TMSE via cross-validation. More specifically, we conduct 10-fold cross-validation and extract $r=3$ PCs sequentially with PCA, predictive PCA and RapPCA on each training dataset, as described in Sections 2 and 3. Our main parameter of interest is $\gamma$, since it controls the predictability-representability trade-off; its behavior will be discussed in this section. Figure 10 in Appendix B.2 provides further details about the impact of $\lambda_{1},\lambda_{2}$ on the solution. Prediction is done by a two-step procedure (see, e.g., Wai et al., 2020), where we first train a random forest model for each of the extracted PCs $\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3}$, and then conduct spatial smoothing on the residuals with thin-plate regression splines (TPRS) (Wood, 2003).
@@ 4.1 Scenario 1: PCs with Equal Contribution
We first consider the setting with 3 (out of 6) well-predictable PCs and 3 PCs mainly consisting of structured and unstructured Gaussian errors. More specifically, we let $f_{l}(X)=X\beta_{l}$ where the entries of $\beta_{l}$ are drawn independently from Uniform[1,1]11[-1,1][ - 1 , 1 ], and $\sigma_{l}^{2}=0$, for $l=1,2,3$; we let $f_{l}=0$ and $\sigma_{l}^{2}=1$ for $l=4,5,6$. $M$ is set to be $\Lambda\tilde{M}$, where $\Lambda$ is diagonal with $\Lambda_{ll}=1/\text{sd}\big{(}\text{PC}^{(l)}\big{)}$ and the entries of $\tilde{M}$ are independent Uniform[1,1]11[-1,1][ - 1 , 1 ]. In other words, we first scale the 6 PCs to have equal variability, and then assign random but overall comparable weights to them.
@@
Comparing the two existing methods, it is expected in this scenario that predictive PCA would outperform classical PCA in prediction without severely compromising approximation accuracy, since the predictable PCs explain a similar amount of variability compared to the unpredictable ones. Consequently, predictive PCA would achieve better overall performance as reflected by TMSE. RapPCA flexibly interpolates between classical and predictive PCA, and is expected to have comparable or improved performance.
@@
Panel A of Figure 1 presents the breakdown of MSPE, MSRE-trn and TMSE for the first PC by $\gamma$, the main tuning parameter for RapPCA, in comparison to PCA and predictive PCA for this scenario. We only present this breakdown for the first PC but not the subsequent ones because different PCA algorithms will have different residuals $Y^{(l)}$ (recall Equations 12 and 3) starting from the second PC, and the PC-specific breakdown of metrics is consequently no longer comparable. We observe that MSPE decreases and training MSRE increases as we increase $\gamma$ for RapPCA, since this imposes a greater penalty on prediction errors in optimization.
@@
Figure 2 compares the prediction MSEs for each PC, as well as the overall TMSE, MSPE and MSRE-trn for different methods. As expected for this scenario, we observe in Panel A that predictive PCA achieves lower prediction errors than PCA, for each PC as well as the overall MSPE. This advantage does not cost a higher approximation gap (MSRE), since the predictable components in the outcome $Y$ explain a similar amount of variability as the unpredictable ones. Our method achieves comparable but improved prediction accuracy than predictive PCA, because it is able to adjust the penalties on the covariate versus spatial effect terms adaptively, especially when there are a large number of spatial basis terms (recall the separate tuning parameters $\lambda_{1}$ and $\lambda_{2}$ in Equation 3), as opposed to predictive PCA which is restricted to a low-dimensional combination of covariates and spatial basis functions (recall the term $Z$ in Equation 2). Consequently, predictive PCA achieves better overall performance (reflected by smaller TMSE) than PCA, while RapPCA further improves the prediction and overall performance due to its flexibility in adaptively capturing the covariate and spatial effects.
@@
Figure 1: Breakdown of MSPE, MSRE-trn (which is MSRE on the training set) and TMSE for the first PC by $\gamma$ arcoss 100 replicates of data, with classical PCA (coded as $\gamma=-1$) and predictive PCA (PredPCA, coded as $\gamma=99$) results presented for reference
@@
Figure 2: Individual prediction MSEs for each PC (left) and overall metrics for all PCs (right) across 100 replicates of data. The optimal tuning parameters $\gamma,\lambda_{1}$ and $\lambda_{2}$ are selected separately for each simulation by minimizing TMSE via cross-validation.
@@ 4.2 Scenario 2: PCs with Unequal Contribution
We now consider a scenario where the predictable PCs explain a higher proportion of variability than the unpredictable PCs. In particular, the simulation setting is similar to Scenario 1, except that the entries of $\tilde{M}$ are first drawn from independent Uniform[1,1]11[-1,1][ - 1 , 1 ] distributions, and then the rows are scaled to have decaying norm (recall that the first 3 PCs are more predictable than the last 3 PCs). In contrast to Scenario 1, we expect in this case that PCA and predictive PCA would have comparable prediction, approximation and overall accuracy, since the well-predictable components of $Y$ coincide with those that explain the greatest amount of variance.
@@ 4.3 Scenario 3: PCs with Non-Linear Mean
Panel C of Figure 1 indicates that the flexibility of handling non-linear effects enables a clear advantage in prediction accuracy for RapPCA compared with both predictive and classical PCA in this scenario, when we impose a large enough penalty $\gamma$ on prediction errors in optimization. It is not surprising that putting an emphasis on prediction leads to some loss in approximation accuracy; however, we could observe from Panel C of Figure 2 that by controlling such balance in a data-adaptive way, RapPCA is able to achieve meaningful improvement in MSPE while maintaining comparable, if not superior, overall performance.
4.4 Remarks on the Predictability-Representability Trade-Off
In all of our simulation studies, we choose the key tuning parameter $\gamma$ of RapPCA data-adaptively by minimizing TMSE in cross-validation. While TMSE is a comprehensive and balanced metric to use for tuning, alternative choices can be made with considerations on the practical use cases.
For example, in health effect studies of air pollution with spatial misalignment between the locations of measured health outcomes and those of measured pollutant concentrations, the accuracy of such spatial extrapolation is arguably more important than the closeness of approximation for the original pollution measurements (Jandarov et al., 2017). In this case, researchers may examine the breakdown of metrics in cross-validation similar to Figure 1 and manually select a value of $\gamma$ leading to satisfying prediction accuracy (MSPE) without significantly compromising MSRE or TMSE. We also note from the curvature of TMSE versus $\gamma$ in these figures that for many practical cases, the overall performance of RapPCA is not too sensitive to the choice of $\gamma$, or the metric(s) used to select it. In other words, manually adjusting the predictability-representability trade-off with RapPCA leads to consistently improved predictability and better, or at least similar, overall accuracy (TMSE) compared with existing methods such as PCA and predictive PCA.
@@ 5 Applications
We illustrate the utility of RapPCA using real datasets in two different domain areas — environmental epidemiology and spatial transcriptomics. The first application of dimension reduction in Section 5.1 ($n=309,\ p=23$) represents the case where practitioners seek to improve predictability (MSPE) and the overall performance (TMSE), and hence RapPCA is able to explicitly optimize these metrics as desired. The second example in Section 5.2 ($n=607,\ p=10,053$) includes the scenario where tasks other than spatial prediction (e.g., clustering) are of interest; we demonstrate that RapPCA, though not directly incorporating the ultimate goal into optimization, is able to capture meaningful biological and spatial information in the extracted PCs by enforcing predictability, leading to desirable model performance.
@@ 5.1 Analysis of Seattle Traffic-Related Air Pollution Data
The performance of RapPCA in comparison with common existing methods, including classical and predictive PCA, is first illustrated with the multivariate traffic-related air pollution (TRAP) data in Seattle (Blanco et al., 2022). The study leverages a mobile monitoring campaign where a vehicle equipped with air monitors repeatedly collected two-minute samples at $n=309$ stationary roadside sites in the greater Seattle area. Repeated samples for the concentrations of 6 types of pollutants were obtained, including ultrafine particles (UFP), black carbon (BC), nitrogen dioxide (NO2), carbon monoxide (CO), carbon dioxide (CO2) and fine particulate matter (PM2.5). Due to data quality considerations, we focus on UFP, BC and NO2 in our analysis. For UFP and BC, the concentrations were measured with multiple instruments corresponding to different measurement ranges or units. In particular, NanoScan instrument with 13 different bin sizes, as well as DiSCmini and PTRAK 8525 (with and without diffusion screen) instruments measured the concentration of UFP in terms of the counts of particles with different size ranges, along with the median size and overall count of particles. Black carbon was assessed by micro-aethalometer (MA200), which measures the concentration of particles with different light absorbing properties, corresponding to 5 different ranges of wavelengths.
@@
The median 2-minute visit concentrations were winsorized at the site level such that concentrations below the 5th and above the 95th quantile for a given site site were set to those thresholds, respectively. This was done to reduce the influence of outliers in the measurements. These winsorized medians were then averaged for each site, leading to $p=23$ annual average measurements of pollutant concentrations in total, including 17 for UFP, 5 for BC, and one for NO2. We use the concentration of NO2 to normalize all other variables except the median size measurements for UFP from the DiSCmini. All variables are centered and scaled before running each PCA algorithm.
@@
We are interested in extrapolating the pollutant measurements to unobserved locations in the same study region. Because of the autocorrelation between these measurements, predicting each of them separately might not always lead to sensible results; instead, dimension reduction is needed before building spatial prediction models. Dimension reduction also allows us to understand the health effects of pollutant mixtures rather than separate pollutants alone. As in Section 4, we compare the performance of RapPCA with classical and predictive PCA, each extracting the top 3 PCs from the 23 measurements of air pollutants. Due to the high dimensionality of geographical covariates, we first ran PCA on these 189 covariates and used the top 15 PCs as predictors for predictive PCA. PCA and RapPCA were applied without this additional step. We examine the overall metrics as well as the individual prediction MSEs for each PC via 10-fold cross-validation, where the same spatial prediction approach as Section 4 is followed, namely, a random forest model followed by spatial smoothing via TPRS on the residuals (Wai et al., 2020).
@@
We then run dimension reduction with each of the PCA algorithms on the whole dataset to assess the spatial distribution of top 3 PC scores, where for RapPCA we select the tuning parameters $\gamma,\lambda_{1}$ and $\lambda_{2}$ that lead to the optimal TMSE. Next, we train spatial prediction models for each of the PC scores, and evaluate these models on a finer grid of 5,040 locations over the study region to obtain smoothed plots of each PC score. Figure 3 visualizes the smoothed PC scores obtained by each method across the study region, and Figure 4 presents the PC loadings reflecting the contribution of each pollutant on the PC scores. We observe similar spatial patterns in the distribution of the PC scores across different methods, except for the south end of the study region for the third PC where PCA identifies a stronger signal than RapPCA and predictive PCA. In particular, all methods highlight regions near the airport and/or around major roads (the dark red area) for the first PC, indicating aircraft and road traffic emissions as a major source of overall pollution level. This is consistent with the large contributions of BC as well as UFP with small or moderate sizes (Zhang et al., 2019; Bendtsen et al., 2021), as reflected by the loadings of UFP (with relatively small particle sizes) and BC in Figure 4.
Roughly categorizing various UFP particle sizes into small, moderate and large groups, one key difference we observe from Figure 4 between RapPCA and the two benchmark algorithms is that RapPCA identifies one size group to be the main contributor to each of the 3 PCs, while the other two groups have negligible loadings: small particles contributing to the top PC, moderately-sized particles to the second PC, and large ones to the third PC. In addition, BC is mainly contributing to the second PC according to RapPCA. This observation is consistent with findings from other studies, where aircraft emissions (primarily composed of small UFPs) and road traffic, especially diesel exhaust (comprising moderately-sized UFPs and BC), were identified as two key sources of traffic-related air pollution (Austin et al., 2021; Nie et al., 2022). In contrast, neither predictive nor classical PCA effectively distinguishes the contributions of UFPs with different sizes or the contribution of BC.
Figure 3: Smoothed PC scores of pollutant concentrations from the Seattle TRAP data
Figure 4: PC loadings for each pollutant. There are 3 types of pollutants, where the suffix, if applicable, represents the properties of, or the instruments used to measure, each pollutant. In particular, the numeric suffix after ufp_ corresponds to the range of sizes for the particles. ufp_ptrak_36 represents PTRAK measurements with diffusion screen, and ufp_ptrak_20 represents the difference between PTRAK measurements without and with diffusion screen. Black carbon (BC) measurements at different wavelengths are shown: blue (bc_blue), green (bc_green), infrared (bc_ir), red (bc_red), and ultraviolet (bc_uv). Ultraviolet measurements were transformed to represent the difference (bc_uv_diff) between ultraviolet and infrared ranges.

### v2

@@ 21      Calculate the $l$th PC score: $\tilde{u}_{l}\leftarrow Y^{(l)}\tilde{v}_{l}$;
For each PC (indexed by $l$), denote the singular value decomposition of $Y^{(l)}$ as $Y^{(l)}=\mathcal{S}^{(l)}{\mathcal{D}^{(l)}}{\mathcal{T}^{(l)}}^{\top}$. Take $Z,P,q$ and the loading $\tilde{v}_{l}$ to be the specified forms in Algorithm 1. Then $(\tilde{v}_{l},\tilde{\alpha}_{l},\tilde{\beta}_{l})$ is the global minimizer of (4), where $\left[\tilde{\alpha}_{l}^{\top}\quad\sqrt{\frac{\lambda_{2}}{\lambda_{1}}}\tilde{\beta}_{l}^{\top}\right]^{\top}=(\gamma Z^{\top}Z+\lambda_{1}P)^{-1}(\gamma Z^{\top}\mathcal{S}^{(l)}{\mathcal{D}^{(l)}}q)$.
The computational complexity of the procedure described above is $O\left((n+m)^{3}+np^{2}\right)$, where we recall that $n$ is the sample size, $p$ is the dimensionality of $Y$, and $m$ is the number of spline basis functions included in $B$. This algorithm is computationally simple in that it obtains the optimal solution with an explicit expression, as opposed to iterative numerical optimization procedures. This algorithm depends on the specification of several quantities. For the hyper-parameters $\gamma^{(l)},\lambda_{1}^{(l)},\lambda_{2}^{(l)}$, one could adopt cross-validation and choose the combination that minimizes TMSE or other metrics (e.g., MSPE/MSRE) depending on the targeted use case, sequentially for each PC. Taking TMSE as an example, for the $l$th PC and for each combination of $\gamma^{(l)},\lambda_{1}^{(l)},\lambda_{2}^{(l)}$, we calculate $\tilde{u}_{l},\tilde{v}_{l}$ on the training set, predict $\hat{u}_{\text{tst}}$ on the validation set, and then examine $\text{TMSE}_{l}=n_{\text{tst}}^{-1}\left\lVert Y^{(l)}_{\text{tst}}-\hat{u}_{\text{tst}}\tilde{v}_{l}^{\top}\right\rVert_{2}^{2}$ to choose the optimal combination minimizing $\text{TMSE}_{l}$. Specification of the dimension $m$ of $B$ is not critical due to the inclusion of penalty on $\beta$ (Wood, 2017). One can start with a large $m$ (e.g., close to the sample size $n$) and the degree of freedom of $\beta$ will be controlled data-adaptively via $\lambda_{2}$.
@@ 4 Simulations
In this section, we compare the performance of RapPCA on simulated data with classical and predictive PCA. In particular, we investigate three scenarios with different trade-offs between predictability and representability resulting from different data generating mechanisms.
@@
$\text{PC}^{(l)}=f_{l}(X)+\epsilon_{l},\ l=1,\ldots,6,$ (5)
@@
$Y_{n\times p}=\text{PC}_{n\times 6}M_{6\times p}+\varepsilon_{n\times p}\,,$
where $M$ is adjusted differently in each scenario to control the contribution from different PCs, and the entries of the noise terms $\varepsilon$ are drawn from independent Normal$(0,0.1)$ distributions.
We examine metrics including TMSE, MSPE and MSRE-trn for 100 replicates of data, in terms of their overall magnitudes, as well as the breakdown by each PC and/or $\gamma$ (the tuning parameter). We also report the individual prediction MSEs for each PC, which is $\text{MSE}_{l}=n_{\text{tst}}^{-1}\lVert\hat{u}_{l}-u^{*}_{l}\rVert_{2}^{2}$, recalling from Section 2.2 that $u^{*}_{l}=Y_{\text{tst}}\tilde{v}_{l}$. We extract $r=3$ PCs sequentially with each of the three candidate algorithms, PCA, predictive PCA and RapPCA. The optimal combination of $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ is selected by minimizing TMSE via 10-fold cross-validation. For each PC score $\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3}$, test set predictions are obtained via a two-step procedure (see, e.g., Wai et al., 2020), where we first train a random forest model, and then conduct spatial smoothing on the residuals with thin-plate regression splines (TPRS) (Wood, 2003). We discuss the role played by the primary tuning parameter $\gamma$ in this section. Figure S3 for RapPCA in Appendix B.2 provides further details about the impact of $\lambda_{1},\lambda_{2}$ on the solution.
@@ 4.1 Scenario 1: PCs with Equal Contribution
We first consider the setting with 3 well-predictable PCs and 3 PCs mainly consisting of structured and unstructured Gaussian errors. More specifically, for the mean model in (5), we let $f_{l}(X)=X\beta_{l}$ where the entries of $\beta_{l}$ are drawn independently from Uniform[1,1]11[-1,1][ - 1 , 1 ], and $\sigma_{l}^{2}=0$, for $l=1,2,3$; we let $f_{l}=0$ and $\sigma_{l}^{2}=1$ for $l=4,5,6$. $M$ is set to be $\Lambda\tilde{M}$, where $\Lambda$ is diagonal with $\Lambda_{ll}=1/\text{sd}\big{(}\text{PC}^{(l)}\big{)}$ and the entries of $\tilde{M}$ are independent Uniform[1,1]11[-1,1][ - 1 , 1 ]. In other words, we first scale the 6 PCs to have equal variability, and then assign random but overall comparable weights to them.
@@
Comparing the two existing methods, it is expected that in this scenario predictive PCA would outperform classical PCA in prediction without severely compromising approximation accuracy, since the predictable PCs explain a similar amount of variability compared to the unpredictable ones. Consequently, predictive PCA would achieve better overall performance as reflected by TMSE. RapPCA flexibly interpolates between classical and predictive PCA, and is expected to have comparable or better performance than predictive PCA.
@@
Panel A in Figure 1 presents the breakdown of MSPE, MSRE-trn and TMSE for the first PC by $\gamma$, in comparison to PCA and predictive PCA for this scenario. We only present this breakdown for the first PC but not the subsequent ones because starting from the second PC, different PCA algorithms will have different residuals $Y^{(l)}$ (recall Equations 12 and 4), and the PC-specific breakdown of metrics is consequently no longer comparable. We observe that MSPE decreases and training MSRE increases as we increase $\gamma$ for RapPCA, since this imposes a greater penalty on prediction errors in optimization.
@@
Figure 2 compares the prediction MSEs for each PC, as well as the overall TMSE, MSPE and MSRE-trn for different methods. As expected for this scenario, we observe in Panel A that predictive PCA achieves lower prediction errors than PCA for each PC, as well as lower overall MSPE. This advantage does not cost a higher approximation gap (MSRE), since the predictable components in the outcome $Y$ explain a similar amount of variability as the unpredictable ones. RapPCA is able to adjust the penalties on the covariate versus spatial effect terms adaptively, especially when there are a large number of spatial basis terms (recall the separate tuning parameters $\lambda_{1}$ and $\lambda_{2}$ in Equation 4), as opposed to predictive PCA which is restricted to a low-dimensional combination of covariates and spatial basis functions (recall the term $Z$ in Equation 2). Consequently, predictive PCA achieves better overall performance (reflected by smaller TMSE) than PCA, while RapPCA further improves the prediction and overall performance due to its flexibility to data-adaptively capture the covariate and spatial effects.
@@
Figure 1: Breakdown of MSPE, MSRE-trn (MSRE on the training set) and TMSE for the first PC by $\gamma$ arcoss 100 replicates of data, with classical PCA (coded as $\gamma=-1$) and predictive PCA (PredPCA, coded as $\gamma=99$) results presented for reference.
@@
Figure 2: Distributions of individual prediction MSEs for each PC (left) and overall metrics for all PCs (right) across 100 replicates of data. The optimal tuning parameters $\gamma,\lambda_{1}$ and $\lambda_{2}$ are selected separately for each simulation by minimizing TMSE via cross-validation.
@@ 4.2 Scenario 2: PCs with Unequal Contribution
We now consider a scenario where the predictable PCs explain a higher proportion of variability than the unpredictable PCs. In particular, the simulation setting is similar to Scenario 1, except that the entries of $\tilde{M}$ are first drawn from independent Uniform[1,1]11[-1,1][ - 1 , 1 ] distributions, and then the rows are scaled to have decaying norm (recall that the first 3 PCs are more predictable than the last 3 PCs). We expect in this case that PCA and predictive PCA would have comparable prediction, approximation and overall accuracy, since the well-predictable components of $Y$ coincide with those that explain the greatest amount of variance.
@@ 4.3 Scenario 3: PCs with Non-Linear Mean
Panel C of Figure 1 indicates that by flexibly handling non-linear effects, RapPCA gains a clear advantage in prediction accuracy compared to both predictive and classical PCA in this scenario, when we impose a large enough penalty $\gamma$ on prediction errors in optimization. It is not surprising that putting an emphasis on prediction leads to some loss in approximation accuracy; however, we observe from Panel C of Figure 2 that by balancing the two objectives data-adaptively, RapPCA achieves meaningful improvement in MSPE while maintaining comparable, if not superior, overall performance.
@@ 5 Applications
We illustrate the utility of RapPCA using real datasets in two different domain areas — environmental epidemiology and spatial transcriptomics. The first application of dimension reduction in Section 5.1 ($n=309,\ p=23$) represents the case where practitioners seek to improve predictability (MSPE) and the overall performance (TMSE), and hence RapPCA is able to explicitly optimize these metrics as desired. The second example in Section 5.2 ($n=607,\ p=10,053$) includes the scenario where tasks other than spatial prediction (e.g., clustering) are of interest; we demonstrate that RapPCA, though not directly incorporating the ultimate goal into its optimization, is able to capture meaningful biological and spatial information in the extracted PCs by enforcing predictability, leading to desirable model performance.
@@ 5.1 Analysis of Seattle Traffic-Related Air Pollution Data
The performance of RapPCA in comparison with common existing methods, including classical and predictive PCA, is first illustrated with the multivariate traffic-related air pollution (TRAP) data in Seattle (Blanco et al., 2022). The study leverages a mobile monitoring campaign where a vehicle equipped with air monitors repeatedly collected two-minute samples at $n=309$ stationary roadside sites in the greater Seattle area. Repeated samples for the concentrations of 6 types of pollutants were obtained, including ultrafine particles (UFP), black carbon (BC), nitrogen dioxide (NO2), carbon monoxide (CO), carbon dioxide (CO2) and fine particulate matter (PM2.5). Following Blanco et al. (2022), we focus on UFP, BC and NO2 in our analysis due to data quality considerations. For UFP and BC, the concentrations were measured with multiple instruments corresponding to different measurement ranges or units. In particular, NanoScan instrument with 13 different bin sizes, as well as DiSCmini and PTRAK 8525 (with and without diffusion screen) instruments measured the concentration of UFP in terms of the counts of particles with different size ranges, along with the median size and overall count of particles. Black carbon was assessed by micro-aethalometer (MA200), which measures the concentration of particles with different light absorbing properties, corresponding to 5 different ranges of wavelengths.
@@
The median 2-minute visit concentrations were winsorized at the site level such that concentrations below the 5th and above the 95th quantile for a given site were set to those thresholds, respectively. This was done to reduce the influence of outliers in the measurements. These winsorized medians were then averaged for each site, leading to $p=23$ annual average measurements of pollutant concentrations in total, including 17 for UFP, 5 for BC, and one for NO2. We use the concentration of NO2 to normalize all other variables except the median size measurements for UFP from the DiSCmini. All variables are centered and scaled before running each PCA algorithm.
@@
We are interested in extrapolating the pollutant measurements to unobserved locations in the same study region. Because of the autocorrelation between these measurements, predicting each of them separately might not always lead to sensible results; instead, dimension reduction is needed before building spatial prediction models. Dimension reduction also allows us to understand the health effects of pollutant mixtures rather than separate pollutants alone. As in Section 4, we compare the performance of RapPCA with classical and predictive PCA, each extracting the top 3 PCs from the 23 measurements of air pollutants. Due to the high dimensionality of geographical covariates, we first ran PCA on these 189 covariates and used the top 15 PCs as predictors for predictive PCA. PCA and RapPCA were applied without this additional step. We examine the overall metrics as well as the individual prediction MSEs for each PC via 10-fold cross-validation following the same spatial prediction approach as Section 4, namely, a random forest model followed by spatial smoothing via TPRS on the residuals (Wai et al., 2020).
@@
We then run dimension reduction with each of the PCA algorithms on the whole dataset to assess the spatial distribution of top 3 PC scores, where for RapPCA we select the tuning parameters $\gamma,\lambda_{1}$ and $\lambda_{2}$ that lead to the optimal TMSE. Next, we train spatial prediction models for each of the PC scores, and evaluate these models on a finer grid of 5,040 locations over the study region to obtain smoothed plots of each PC score. Figure S4 in Appendix B.3 visualizes the smoothed, finer-grain PC scores obtained by each method across the study region. Figure 3 presents the PC loadings reflecting the contribution of each pollutant on the PC scores. Roughly categorizing various UFP particle sizes into small, moderate and large groups, one key difference we observe from Figure 3 between RapPCA and the two benchmark algorithms is that RapPCA identifies one size group to be the main contributor to each of the 3 PCs, while the other two groups have negligible loadings: small particles contributing to the top PC, moderately-sized particles to the second PC, and large ones to the third PC. In addition, BC is mainly contributing to the second PC according to RapPCA. This observation is consistent with findings from other studies, where aircraft emissions (primarily composed of small UFPs) and road traffic, especially diesel exhaust (comprising moderately-sized UFPs and BC), were identified as two key sources of traffic-related air pollution (Austin et al., 2021; Nie et al., 2022). These findings reflect the benefits in downstream model interpretability brought by dimension reduction that well preserves scientific and spatial information. In contrast, neither predictive nor classical PCA effectively distinguishes the contributions of UFPs with different sizes or the contribution of BC.
Figure 3: PC loadings for each of the top 3 PCs obtained by PCA, predictive PCA and RapPCA for each pollutant. There are 3 types of pollutants, UFP, BC and NO2. The suffix, if applicable, represents the properties of, or the instruments used to measure, each pollutant. In particular, the numeric suffix after ufp_ corresponds to the range of sizes for the particles. ufp_ptrak_36 represents PTRAK measurements with diffusion screen, and ufp_ptrak_20 represents the difference between PTRAK measurements without and with diffusion screen. Black carbon (BC) measurements corresponding to different wavelengths are shown: blue (bc_blue), green (bc_green), infrared (bc_ir), red (bc_red), and ultraviolet (bc_uv). Ultraviolet measurements were transformed to represent the difference (bc_uv_diff) between ultraviolet and infrared ranges.